Q4 On mountainous downhill roads, escape routes are sometimes placed on the side of the road for t Assuming a constant upward slope of 26, calculate the horizontal that slowed from 110 km to rest in 7.0 seconds. (see proble Ch 3, page 72, #53). rucks whose brakes might fail. and vertical components of the acceleration of a truck

please help me answer this question and also please write clearly so I can understand thank you so much also please show all steps thank you so much
Answer #1

According to the question provided we know that

theta = 26 deg d = 110 km/h = 30.555 m/s t =7.00 seconds ===provided values

a = 0 - 30.5555 m/s / 7.00 s = 4.365079 m/s^2

Part a) 4.365079 m/s^2 * cos 26 deg = 3.9233 m/s^2 ==========ANSWER)

Part b) 4.365079 m/s^2 * sin 26 deg = 1.913524 m/s^2 =========ANSWER)